# LexC-Gen: Generating Data for Extremely Low-Resource Languages with Large Language Models and Bilingual Lexicons

### v1

@@ Abstract
Data scarcity in low-resource languages can be addressed with word-to-word translations from labeled task data in high-resource languages using bilingual lexicons. However, bilingual lexicons often have limited lexical overlap with task data, which results in poor translation coverage and lexicon utilization. We propose lexicon-conditioned data generation (LexC-Gen), a method that generates low-resource-language classification task data at scale. Specifically, LexC-Gen first uses high-resource-language words from bilingual lexicons to generate lexicon-compatible task data, and then it translates them into low-resource languages with bilingual lexicons via word translation. Across 17 extremely low-resource languages, LexC-Gen generated data is competitive with expert-translated gold data, and yields on average 5.6 and 8.9 points improvement over existing lexicon-based word translation methods on sentiment analysis and topic classification tasks respectively. We show that conditioning on bilingual lexicons is the key component of LexC-Gen. LexC-Gen is also practical—it only needs a single GPU to generate data at scale. It works well with open-access LLMs, and its cost is one-fifth of the cost of GPT4-based multilingual data generation.
@@
[…]
@@ 1 Introduction
Labeled data are virtually non-existent for extremely low-resource languages (Joshi et al., 2020), which creates significant disparities in NLP advancements relative to high-resource languages (Mabokela et al., 2022; Robinson et al., 2023; Bang et al., 2023; Yong et al., 2023a; Üstün et al., 2024). One solution to overcome the data bottleneck is to use bilingual lexicons, which cover more than 5000 languages around the world Wang et al. (2022); Koto et al. (2024). Bilingual lexicons, often the first product of language documentation (Meara, 1993; Schreuder and Weltens, 1993; Kroll and Ma, 2017), are dictionaries that map words from one language to their translations in another language. With bilingual lexicons, one can obtain data in low-resource languages by translating labeled task data from high-resource languages through word-for-word substitution.
@@
Figure 1: We observe data-lexicon mismatch (i.e., low lexical overlap) between existing task data and bilingual lexicons (Figure 0(a)). LexC-Gen addresses the issue by maximizing the lexical overlap. Its generated data have more words translated (i.e., higher word translation coverage) and cover more low-resource-language words in bilingual lexicon (i.e., higher lexicon utilization rate) (Figure 0(b)).
@@
While prior work has demonstrated the effectiveness of bilingual lexicons in augmenting data and improving downstream task performance (Wang et al., 2022; Jones et al., 2023, inter alia), we observe that often the words in the existing task data—readily available labeled data in high-resource languages for a target task, e.g., sentiment analysis or topic classification—have low lexical overlap with the words in the task-agnostic bilingual lexicons, as shown in Figure 1. This data-lexicon mismatch creates two problems: (1) many high-resource-language word tokens remain untranslated and (2) many words in the bilingual lexicon, which possibly contain useful semantic information for downstream tasks, are missing from the translated dataset.
In this work, we introduce LexC-Gen,111pronounced as lek-see-jen which is a lexicon-conditioned data generation method, to mitigate data-lexicon mismatch and generate task training data for extremely low-resource languages. Specifically, we train LLMs to generate high-resource-language task data using words from bilingual lexicons, so the data have a higher lexical overlap with the lexicons. This results in better word translation coverage and lexicon utilization rate (Figure 1). Then we translate the generated data into low-resource languages using bilingual lexicons. We also propose a quality-control method that checks for input-label consistency to filter out poor-quality generated data.
@@
LexC-Gen’s data generation process is cost-effective. Specifically, it takes only a single V100 GPU and less than 36 hours to complete data generation with LexC-Gen at scale—this costs less than \$100 with Google Cloud services. This is 20% of the cost of GPT4-based method for multilingual data generation Whitehouse et al. (2023).
LexC-Gen also works well with LLMs with permissive licenses such as BLOOMZ Muennighoff et al. (2023). Therefore, the generated cross-lingual training data can be made open access for further research and building systems for extremely low-resource languages, which benefits multilingual NLP progress for these data-scarce languages.
@@
    We present LexC-Gen, a lexicon-conditioned generation method that conditions LLMs on bilingual lexicons to generate low-resource-language task data at scale.
@@
    We performed extensive ablation study on LexC-Gen. We show that simply scaling up generated task data is insufficient. Lexicon-conditioning is necessary to maximize lexical overlap between task data and bilingual lexicons.
@@ 2 Related Work
Lexicon-based augmentation creates data for low-resource languages by swapping words in high-resource-language data with their dictionary word translations in bilingual lexicons. This is useful for low-resource languages that cannot be readily translated by translation models/APIs with limited language coverage. Prior work has demonstrated their effectiveness across a wide range of NLP tasks, such as machine translation (Streiter and Iomdin, 2000; Ramesh and Sankaranarayanan, 2018; Thompson et al., 2019; Kumar et al., 2022; Jones et al., 2023), sequence labeling (Scherrer and Sagot, 2013; Mayhew et al., 2017; Wang et al., 2022), sentiment classification (Rasooli et al., 2018; Ali et al., 2021; Mohammed and Prasad, 2023), and topic classification Song et al. (2019). However, many lexicon-based data augmentation strategies for semantic tasks in low-resource languages rely on domain-specific lexicons Das and Bandyopadhyay (2010); Buechel et al. (2016); Ali et al. (2021); Mohammed and Prasad (2023); Koto et al. (2024), and performance-wise they still fall short of gold training data collected in the target low-resource language Rasooli et al. (2018); Koto et al. (2024). Our method LexC-Gen not only works with domain-agnostic bilingual lexicons, but also demonstrates competitive performance with gold training data on sentiment analysis and topic classification tasks across many extremely low-resource languages.
@@
Zero-shot prompting
@@
Table 1: Sentiment analysis accuracy on 7 Indonesian extremely low-resource local languages in the NusaX dataset Winata et al. (2023b). (T) indicates word-translated data, and (en) refers to the existing task data in English. The terms -1K, -10K and -100K refer to the size of synthetic training data generated by our proposed LexC-Gen before filtering. Column “#data” refers to the actual training data size. We shade the gold translations where the mBERT classifier is trained on expert-translated data in low-resource languages created with native speakers’ translations. We bold the best result and underline the second-best. We report results averaged over 5 seeds of classifier finetuning.
@@ 3.4 Word-to-Word Translation into Low-Resource Languages
Zero-shot prompting
@@ 4.2 LexC-Gen Instantiation
We provide further training and inference details of LexC-Gen in Appendix C. We also showcase examples of the generated data in Appendix E.
@@ 4.3 Baselines
We compare LexC-Gen against (1) zero-shot prompting with BLOOMZ-7.1B and GPT-4; [^3] We used the most recent gpt-4-0125-turbo. See Appendix F for more details. (2) cross-lingual zero-shot transfer where mBERT is finetuned on English training data and evaluated on low-resource-language test data; (3) word translation (Wang et al., 2022) where mBERT is finetuned on the data that are translated from the English training data via word-substitution with the same bilingual lexicon Gatitos (Jones et al., 2023); (4) gold translations where mBERT is finetuned on expert-translated task training data in the target low-resource language (see Section 4.1)
@@ 5 Results and Analysis
5.1 LexC-Gen outperforms all baselines
LexC-Gen outperforms all baselines in both sentiment analysis (Section 2) and topic classification tasks (Section 3.4). In sentiment analysis, finetuning classifiers on the mixture of LexC-Gen-100K (100K generated data instances that are filtered down to around 37K instances) and existing English task data improves over the cross-lingual zero-shot baseline by 15.2 percentage points and word translation baseline by 6.6 points. In topic classification, LexC-Gen-100K yields improvement of 18.3 points over the cross-lingual zero-shot baseline and 8.9 points over the word translation baseline.
@@ 5.2 LexC-Gen is competitive with gold translations
Section 2 and Section 3.4 show that finetuning classifiers on LexC-Gen-100K generated cross-lingual data is competitive with training on expert-translated data for many low-resource languages. Our findings also generalize to larger task classifiers, such as XLMR-base and XLMR-large (Conneau et al., 2020) (see Figure 10 in Appendix J). Our result is surprising because LexC-Gen generated data still use English syntax with SVO word order. Yet, LexC-Gen still works for languages with different word orders, such as Balinese (ban) and Mizo (lus) with OSV word order and Toba batak (bbc) with VOS word order.
@@
Figure 4: Ablation study of lexicon-conditioning in LexC-Gen-100K on sentiment analysis. “Gen w/o filter” generates data without using words from lexicons and without input-label consistency filtering. “Gen” is “Gen w/o filter” but with filtering. We plot the accuracy difference against finetuning with gold translations (green dotted line). We control the training data size for “Gen w/o filter” and “Gen” to be the same as LexC-Gen-100K and include (Wang et al., 2022) as baseline.
@@
5.3 Mixing in English task data helps for small-scale translated data
@@ 5.4 Lexicon-conditioning is crucial for strong task performance
Figure 4 shows that using words from lexicons to generate task data (i.e., lexicon-conditioning) is necessary for matching gold translations performance. Ablating away lexicon-conditioning and quality control (“Gen w/o filter”) has the worst performance— it even underperforms the word translation baseline (Wang et al., 2022) on 500 existing task data samples for sentiment analysis. Even with quality control from Section 3.4, scaling data generation without lexicon conditioning (“Gen”) still performs worse than LexC-Gen-100K. This is due to low lexical overlap between the data and bilingual lexicons. “Gen” data have poorer lexicon utilization rate, as it only covers 62.5% of low-resource-language words in the bilingual lexicon. In contrast, LexC-Gen-100K covers 92.8% words. We refer our readers to Appendix G for further details of our ablation study.
@@ 5.6 Quality control reduces training data size and boosts performance
Quality control with a classifier trained on existing task data is effective for LexC-Gen, but not for label distillation in Wang et al.’s (2022) word translation baseline (Section 2 and Section 3.4). There are two possible reasons. First, label distillation uses the classifier trained on high-resource-language data to relabel translated data in low-resource languages. This cross-lingual transfer may introduce errors in the classifier’s predictions, as opposed to LexC-Gen’s relabeling in the same high-resource language. Second, LexC-Gen offers stricter quality control by discarding all instances with mismatched labels between the classifier and LLMs, thus improving task performance (see Figure 12 in Appendix M).
5.7 LLMs are better used for generating data than zero-shot prompting
Zero-shot prompting with BLOOMZ-7.1B is the weakest baseline (Section 2 and Section 3.4), but using it in LexC-Gen to generate task data to finetune smaller task classifiers makes it achieve the state-of-the-art performance and match gold translations. Our results suggest that for extremely low-resource languages it is best to leverage LLMs (including GPT-4 as shown in Appendix H) to generate training data at scale instead of using them out-of-the-box by zero-shot prompting. [^5] Even few-shot prompting GPT-4 cannot close the performance gap with gold translations. See Appendix K. This finding aligns with recent work that shows that finetuning LLMs with their self-generated task data improves downstream task performance Wang et al. (2023).
@@ 6 Discussion
Effectiveness of bilingual lexicons
Our work shows that task-agnostic bilingual lexicons like Gatitos (Jones et al., 2023) contain sufficient semantic information for sentiment analysis and topic classification in extremely low-resource languages. However, it requires a high degree of lexical overlap between task data and lexicon to include the information in the translated data (Figure 0(a)). We also found that lexicon size and quality are important. Using Gatitos lexicons (Jones et al., 2023) for LexC-Gen outperforms using Panlex (Kamholz et al., 2014) because the former contains more entries and is higher in quality for extremely low-resource languages (see Appendix I).
Addressing training data bottleneck
LexC-Gen directly addresses the labeled data scarcity problem of extremely low-resource languages (Joshi et al., 2020). Since it is challenging to construct high-quality task-specific lexicons or to collect labeled task data that align well with bilingual lexicons, LexC-Gen serves a practical solution that automatically generates task training data at scale for low-resource languages.
Low LLM training cost
LexC-Gen relies on the CTG-trained LLM that follows the prompt instruction of generating task data using a set of given words. Our CTG training only depends on high-resource-language task data, and is independent of low-resource languages and bilingual lexicons. In other words, once an LLM is CTG-trained, researchers can reuse it with different bilingual lexicons to generate data for various low-resource languages on the same task without retraining, thus incurring minimal LLM training cost.
Low LLM generation cost
LexC-Gen benefits low-resource language NLP practitioners who often face resource constraint. LexC-Gen only takes less than 36 hours in total to complete both CTG training and 100K data samples generation with a 7-billion-parameter LLM on a single Tesla V100 GPU. Overall, it costs less than \$100.666A Tesla V100 GPU costs around \$2.48 per hour for on-demand VMs on Google Cloud. This is one-fifth of the cost of using state-of-the-art GPT-4-based multilingual data generation method Whitehouse et al. (2023) (see Appendix H).
Distribution-friendly synthetic data
We show that LexC-Gen works well with open-access models with permissive licenses such as BLOOMZ. Therefore, its generated data are usable for proprietary or public research. The data can also be distributed for broader multilingual applications.

### v2

@@ Abstract
Data scarcity in low-resource languages can be addressed with word-to-word translations from labeled task data in high-resource languages using bilingual lexicons. However, bilingual lexicons often have limited lexical overlap with task data, which results in poor translation coverage and lexicon utilization. We propose lexicon-conditioned data generation (LexC-Gen), a method that generates low-resource-language classification task data at scale. Specifically, LexC-Gen first uses high-resource-language words from bilingual lexicons to generate lexicon-compatible task data, and then it translates them into low-resource languages with bilingual lexicons via word translation. Across 17 extremely low-resource languages, LexC-Gen generated data is competitive with expert-translated gold data, and yields on average 5.6 and 8.9 points improvement over existing lexicon-based word translation methods on sentiment analysis and topic classification tasks respectively. Through ablation study, we show that conditioning on bilingual lexicons is the key component of LexC-Gen. LexC-Gen serves as a potential solution to close the performance gap between open-source multilingual models such as BLOOMZ and state-of-the-art commercial models like GPT-4o on low-resource-language tasks.
@@
Zheng-Xin Yong1  Cristina Menghini2  Stephen H. Bach1 1Department of Computer Science, Brown University 2Data Science Institute, Brown University {contact.yong,cristina_menghini,stephen_bach}@brown.edu
@@ 1 Introduction
Extremely low-resource languages do not have any labeled data and are thereby considered the “Left-Behinds” in NLP language technology development (Joshi et al., 2020; Mabokela et al., 2022; Robinson et al., 2023). Nonetheless, many of them have bilingual lexicons resources, which are usually the first product of language documentation (Meara, 1993; Schreuder and Weltens, 1993; Kroll and Ma, 2017). Bilingual lexicons are dictionaries that map words from one language to their translations in another languages, and they cover more than 5000 languages around the world Wang et al. (2022); Koto et al. (2024).
@@
Figure 1: We observe data-lexicon mismatch (i.e., low lexical overlap) between existing task data and bilingual lexicons (Figure 1(a)). LexC-Gen addresses the issue by generating data using words from lexicons so the data will have more words translated (i.e., higher word translation coverage) and higher lexicon utilization rate (Figure 1(b)).
@@
Previous work uses bilingual lexicons to directly translate labeled data from high-resource languages to low-resource languages through word-for-word substitution (Wang et al., 2022; Jones et al., 2023, inter alia). However, we argue that it is ineffective because of data-lexicon mismatch. Often, the words in the existing task data—readily available labeled data in high-resource languages for a target task, e.g., sentiment analysis or topic classification—have low lexical overlap with the words in the task-agnostic bilingual lexicons, as shown in Figure 1. This mismatch not only results in many words remain untranslated, but also causes entries in the bilingual lexicon, which possibly contain useful semantic information for downstream tasks, missing from the translated dataset.
In this work, we introduce LexC-Gen,111pronounced as lek-see-jen which is a lexicon-conditioned data generation method, to mitigate data-lexicon mismatch through synthetic data generation. Specifically, we train LLMs to generate task data using words from bilingual lexicons, so the data have a higher lexical overlap with the lexicons. This results in better word translation coverage and lexicon utilization rate (Figure 1). We also propose a quality-control method that checks for input-label consistency to filter out poor-quality generated data.
@@
Finally, we discuss how LexC-Gen helps close the performance gap of open-source LLMs in low-resource-language tasks. We show that instead of zero-shot or few-shot prompting, it is better to use them to generate training data with LexC-Gen. The data generation process is cost-effective, and the permissive nature of the models allows generated data to be made open access for further research and building systems for extremely low-resource languages, which benefits multilingual NLP progress for these data-scarce languages.
@@
    We present LexC-Gen, a method that conditions LLMs on bilingual lexicons to generate low-resource-language task data to address data-lexicon mismatch problem.
@@
    Our extensive ablation study on LexC-Gen shows that simply scaling up generated task data is insufficient. Lexicon-conditioning is necessary to maximize lexical overlap between task data and bilingual lexicons.
@@ 2 Related Work
Lexicon-based augmentation creates data for low-resource languages by swapping words in high-resource-language data with their dictionary word translations in bilingual lexicons. This is useful for low-resource languages that cannot be readily translated by translation models/APIs with limited language coverage. Prior work has demonstrated their effectiveness across a wide range of NLP tasks, such as machine translation (Streiter and Iomdin, 2000; Ramesh and Sankaranarayanan, 2018; Thompson et al., 2019; Kumar et al., 2022; Jones et al., 2023), sequence labeling (Scherrer and Sagot, 2013; Mayhew et al., 2017; Wang et al., 2022), sentiment classification (Rasooli et al., 2018; Ali et al., 2021; Mohammed and Prasad, 2023), and topic classification Song et al. (2019). However, many lexicon-based data augmentation strategies for semantic tasks in low-resource languages rely on domain-specific lexicons Das and Bandyopadhyay (2010); Buechel et al. (2016); Mohammed and Prasad (2023); Koto et al. (2024), and performance-wise they still fall short of gold training data collected in the target low-resource language Rasooli et al. (2018); Koto et al. (2024). Our method LexC-Gen not only works with domain-agnostic bilingual lexicons, but also demonstrates competitive performance with gold training data on sentiment analysis and topic classification tasks across many low-resource languages.
@@
Zero-shot/Few-shot prompting
@@
Table 1: Sentiment analysis accuracy on 7 Indonesian extremely low-resource local languages in the NusaX dataset Winata et al. (2023b). (T) indicates word-translated data, and (en) refers to the existing task data in English. The terms -1K, -10K and -100K refer to the size of training data generated by LexC-Gen before filtering. We bold the best result and underline the second-best. We report results averaged over 5 seeds of classifier finetuning.
@@ 3.4 Word-to-Word Translation into Low-Resource Languages
Zero-/Few-shot prompting
@@ 4.2 LexC-Gen Instantiation
We provide further training and inference details of LexC-Gen in Appendix C. We also showcase examples of the generated data for sentiment analysis in Table 3 and for topic classification in Table 6.
Generated Text Sentiment
ulon ’m reusam leumeeh ngeun hek , ulee sikula papeun tuleh member . Hike trails , ta’jub jamek let man keun keu lon . Negative
Please , peutamah nyan pre uteun handbook jadwal keulayi keu umum ureung umum , nyan ’s jareung hadiah lam nyan areusip Neutral
Wonderful , trang ngeun mangat , superior guna , tajam ngeun carong , ngeun nyan barang nakeuh superb . ulon nasihat meujuang toke ’s ho jak keu nyan , nyan ’s saboh konfiden peuningkat . Positive
@@ 4.3 Baselines
We compare LexC-Gen against (1) zero-shot prompting with BLOOMZ-7.1B, Aya-101-13B (Üstün et al., 2024) and GPT-4o; [^3] We used the latest version gpt-4o-2024-05-13. See Appendix E for more details. (2) few-shot prompting with Aya-101-13B using five in-context learning examples; (3) cross-lingual zero-shot transfer where mBERT is finetuned on English training data and evaluated on low-resource-language test data; (4) word translation (Wang et al., 2022) where mBERT is finetuned on the data that are translated from the English training data via word-substitution with the same bilingual lexicon Gatitos (Jones et al., 2023); (5) gold translations where mBERT is finetuned on expert-translated task training data in the target low-resource language (see Section 4.1)
@@ 5 Results and Analysis
5.1 LexC-Gen improves over open-source LLMs and direct word translation
LexC-Gen outperforms prompting open-source models, such as BLOOMZ and Aya-101, and word translation baselines in both sentiment analysis (Section 2) and topic classification tasks (Section 3.4). In sentiment analysis, finetuning classifiers on the mixture of LexC-Gen-100K (100K generated data instances that are filtered down to around 37K instances) and existing English task data improves over the cross-lingual zero-shot baseline by 15.2 percentage points and word translation baseline by 6.6 points. In topic classification, LexC-Gen-100K yields improvement of 18.3 points over the cross-lingual zero-shot baseline and 8.9 points over the word translation baseline. The accuracy gain from adding existing English data reduces from LexC-Gen-1K to LexC-Gen-100K because the English data are dominated by the substantially larger size of generated data (see more discussion in Appendix I).
While the commercially available model GPT-4o yields the best performance—even surpassing classifiers trained on gold data—it is unclear whether the evaluation data has been seen during training. Furthermore, the release of GPT-4o is subsequent to our work. In contrast, our evaluation tasks of NusaX and SIB-200 are not part of the training of open-source models BLOOMZ-7.1B and Aya-101-13B (Workshop et al., 2022; Singh et al., 2024; Üstün et al., 2024). Our results reveal the performance gap in these open-source models. For instance, zero-shot prompting with BLOOMZ-7.1B is the weakest baseline (Section 2 and Section 3.4). However, using it in LexC-Gen to generate task data (i.e., LexC-Gen-100K) can achieve state-of-the-art performance. Our results suggest that, for applying open-source LLMs to low-resource-language tasks, it is best to leverage them to generate training data at scale instead of prompting them directly in zero-shot or few-shot settings.
@@ 5.1 LexC-Gen improves over open-source LLMs and direct word translation
Figure 4: Ablation study of lexicon-conditioning in LexC-Gen-100K on sentiment analysis. The plot shows that accuracy difference against finetuning with gold translations (green dotted line).
@@ 5.2 LexC-Gen is competitive with gold translations
Section 2 and Section 3.4 show that finetuning classifiers on LexC-Gen-100K generated cross-lingual data is competitive with training on expert-translated data for many low-resource languages. Our findings also generalize to larger task classifiers, such as XLMR-base and XLMR-large (Conneau et al., 2020) (see Figure 7 in Section 5.6). Our result is surprising because LexC-Gen generated data still use English syntax with SVO word order. Yet, LexC-Gen still works for languages with different word orders, such as Balinese (ban) and Mizo (lus) with OSV word order and Toba batak (bbc) with VOS word order.
@@ 5.3 Lexicon-conditioning is crucial for strong task performance
Figure 4 shows that using words from lexicons to generate task data (i.e., lexicon-conditioning) is necessary for matching gold translations performance. Ablating away lexicon-conditioning and quality control (“Gen w/o filter”) has the worst performance— it even underperforms the word translation baseline (Wang et al., 2022) on 500 existing task data samples for sentiment analysis. Even with quality control from Section 3.4, scaling data generation without lexicon conditioning (“Gen”) still performs worse than LexC-Gen-100K. This is due to low lexical overlap between the data and bilingual lexicons. “Gen” data have poorer lexicon utilization rate, as it only covers 62.5% of low-resource-language words in the bilingual lexicon. In contrast, LexC-Gen-100K covers 92.8% words. We refer our readers to Appendix F for further details of our ablation study.
@@ 5.5 Quality control reduces training data size and boosts performance
Quality control with a classifier trained on existing task data is effective for LexC-Gen, but not for label distillation in Wang et al.’s (2022) word translation baseline (Section 2 and Section 3.4). There are two possible reasons. First, label distillation uses the classifier trained on high-resource-language data to relabel translated data in low-resource languages. This cross-lingual transfer may introduce errors in the classifier’s predictions, as opposed to LexC-Gen’s relabeling in the same high-resource language. Second, LexC-Gen offers stricter quality control by discarding all instances with mismatched labels between the classifier and LLMs, thus improving task performance (see Figure 11 in Appendix H).
5.6 LexC-Gen generalizes to larger classifiers
Figure 7 breaks down the LexC-Gen generated data size required for task classifiers of different sizes—mBERT (Devlin et al., 2019) has 172 million parameters, XLMR-base (Conneau et al., 2020) has 270 million parameters, and XLMR-large has 550 million parameters—to match gold translations performance. First, we observe that larger task classifier size requires less data to achieve same accuracy. For instance, XLMR-large already exceeds accuracy of 70 points with 5K LexC-Gen data but mBERT requires 35K LexC-Gen data to reach the same accuracy.
However, to match gold performance, the amount of LexC-Gen data does not correlate with the size of classifiers. We find that XLMR-base matches gold performance at around 15K, as opposed to mBERT at around 35K, but XLMR-large requires around 10K more LexC-Gen data than XLMR-base to be as competitive as gold translations.
5.7 Choice of lexicons matters
We compare the usage of Gatitos (Jones et al., 2023) against Panlex (Kamholz et al., 2014) as the lexicons for LexC-Gen. Figure 8 shows that translating LexC-Gen generated data with Gatitos outperforms translating with Panlex on NusaX. One reason is that Panlex has a smaller lexicon size than Gatitos, as for all seven languages in NusaX, Panlex only has around 840 entries, but Gatitos has around 4271 entries. Therefore, LexC-Gen data have a poorer word translation coverage with Panlex. Furthermore, while the data source of Gatitos is not detailed by Jones et al. (2023), the authors described that Gatitos entries are manually reviewed. In other words, the word translations with Gatitos are of higher quality.
@@ 6 Discussion
Application in resource-scarce scenarios
LexC-Gen addresses the resource-scarce scenarios faced by extremely low-resource languages that lack labeled data. We show that we can leverage their existing-yet-scarce lexical resources such as Gatitos (Jones et al., 2023), which only contains around a few thousand of lexicon entries for common words or phrases, to generate labeled task data. Furthermore, LexC-Gen is a practical solution to data-lexicon mismatch problem, as it does not require linguists to build task-specific lexicons, such as multilingual sentiment lexicons (Chen and Skiena, 2014), or practitioners to collect labeled task data in low-resource languages.
Open-source models
Current open-source models like BLOOMZ (Muennighoff et al., 2023) and Aya-101 (Üstün et al., 2024) fail to close the performance gap against GPT-4o and gold translation baseline. Our work bridges the gap—we show that using them to generate training data improves performance over direct zero-shot or few-shot prompting and can match training classifiers on human-labeled data. Furthermore, due to the permissive nature of the models, their generated data can be used for proprietary or public research for broader multilingual applications.
Effectiveness of lexicon-conditioned generation
We shows that task-agnostic bilingual lexicons like Gatitos (Jones et al., 2023) al contain sufficient semantic information for sentiment analysis and topic classification in extremely low-resource languages. However, it requires a high degree of lexical overlap between task data and lexicon to include the information in the translated data (Figure 1(a)). We also found that lexicon size and quality are important. Using Gatitos lexicons (Jones et al., 2023) for LexC-Gen outperforms using Panlex (Kamholz et al., 2014) because the former contains more entries and is higher in quality.
LexC-Gen differs from prior work on lexically constrained text generation (Hokamp and Liu, 2017; Post and Vilar, 2018; Hu et al., 2019). We introduce an additional step of CTG training so LLMs can learn to generate natural text that both maximizes lexicon usage and aligns with class labels. This step allows LexC-Gen to outperform lexically constrained decoding (see Appendix G).
Cost-effectiveness
LexC-Gen relies on the CTG-trained LLM that follows the prompt instruction of generating task data using a set of given words. Our CTG training of open-source LLMs only depends on high-resource-language task data, and is independent of low-resource languages and bilingual lexicons. In other words, once an LLM is CTG-trained, researchers can reuse it with different bilingual lexicons to generate data for various low-resource languages on the same task without retraining. Furthermore, LexC-Gen only takes less than a day to generate 100K data samples on one V100 GPU.
Bilingual lexicon induction (BLI)
We analyze the generated data and discover that on average 34% of the high-resource-language words cannot be found in Gatitos and thus cannot be translated. This leaves room for improvement with BLI to expand the word coverage of bilingual lexicons (Nasution et al., 2016; Irvine and Callison-Burch, 2017; Bafna et al., 2024), so that more words can be translated into low-resource languages. Nonetheless, given that LexC-Gen already matches gold performance, we leave enhancing LexC-Gen with BLI to future work.